How do you solve this linear equation using substitution?
X-y=11
5x-4y=5

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

How do you solve this linear equation using substitution?
X-y=11
5x-4y=5
----
x = y+11
----
Substitute for "x" and solve for "y":
5(y+11)-4y = 5
5y+55-4y = 5
y = -50
-----
Solve for "x":
x = y + 11
x = -50 + 11 = -39
----------------------
Cheers,
Stan H.
=======================

Answer by mananth(16949)

5 x -4 y = 5 .............1
1 y = 11 -1 x ---------2
/ 1 =
y = 11 -1 x
Substitute y in (1)
5 x -4 ( 11 -1 x ) =
5 x -44 + 4 x = 5
5 x + 4 x = 5 44
9 x = 49
/ 9
x= 49/9
Plug the value of x in (1)
5 x -4 y = 5
5 * 49/9 -4 y = 5
245/9 -4 y = 5
-4 y = 5-(249/9)
-4y=-204/9
/ -4
y= 51/9
m.ananth@hotmail.ca